I am supposed to graph the line 3y=-2 but these problems confuse me, could you help
I am supposed to get the two points that make up this line

Answer by mananth(16946)

3y = -2
compare with the equation y = mx+b
x=0
3y=-2
/3
y=-2/3
So the point is (0,-2/3)
There is no x intercept so the line is parallel to the x axis.